OpenStudy (anonymous):

solve for x: 3=x^3+x^2

OpenStudy (anonymous):

3=x^3+x^2 3=X^5 - You can add the expoenents because they have the same base X \[\sqrt[5]{3}=X\]

OpenStudy (anonymous):

I thought you could only do that if they were multiplied